The ledger of Marin Inc. at the end of the current year shows Accounts Receivable $86,000; Credit Sales $800,000; and Sales Returns and Allowances $44,000.

(a) If Marin uses the direct write-off method to account for uncollectible accounts, journalize the adjusting entry at December 31, assuming Marin determines that Matisse’s $700 balance is uncollectible.
(b) If Allowance for Doubtful Accounts has a credit balance of $1,300 in the trial balance, journalize the adjusting entry at December 31, assuming bad debts are expected to be 10% of accounts receivable.
(c) If Allowance for Doubtful Accounts has a debit balance of $450 in the trial balance, journalize the adjusting entry at December 31, assuming bad debts are expected to be 8% of accounts receivable.


Prepare journal entries to record the above transactions. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

No.

Date

Account Titles and Explanation

Debit

Credit

(a)

Dec. 31

enter an account title to record the first transaction

Entry field with correct answer

enter a debit amount

Entry field with correct answer

enter a credit amount

Entry field with correct answer

enter an account title to record the first transaction

Entry field with correct answer

enter a debit amount

Entry field with correct answer

enter a credit amount

Entry field with correct answer

(b)

Dec. 31

enter an account title to record the second transaction

Entry field with correct answer

enter a debit amount

Entry field with incorrect answer

enter a credit amount

Entry field with correct answer

enter an account title to record the second transaction

Entry field with correct answer

enter a debit amount

Entry field with correct answer

enter a credit amount

Entry field with incorrect answer

(c)

Dec. 31

enter an account title to record the third transaction

Entry field with correct answer

enter a debit amount

Entry field with incorrect answer

enter a credit amount

Entry field with correct answer

enter an account title to record the third transaction

Entry field with correct answer

enter a debit amount

Entry field with correct answer

enter a credit amount

Entry field with incorrect answer
Answer #1
date particulars Debit Credit
Dec-31 Bad debt expense 700
Accounts receivable 700
(being accounts receivable written off)
Dec-31 bad debt expense 82900 ( Accounts receivable= 86000+800000-44000=842000)
Allowance for Doubtful Accounts 82900 Bad debt estimate= 842000 x 10% = 84200
( Being adjustment entry passed ) balance in allowance for doubtful accounts= 1300 credit
adjusting entry= 84200-1300=82900
Dec-31 bad debt expense 67810 bad debt estimate= 842000 x 8%=67360
Allowance for Doubtful Accounts 67810 balance in allowance for doubtful accounts= 450 debit
( Being adjustment entry passed ) adjusting entry=67360+450 = 67810